# Robot-mediated physical Human-Human Interaction in Neurorehabilitation: a position paper

### v1

@@ V-F Remote Therapy
Telerehabilitation (rehabilitation delivered remotely) has proven effective in improving motor function, cognition, balance, speech, and communication skills [95]. Studies have shown that telerehabilitation can yield outcomes comparable, or even superior, to conventional rehabilitation [96]. A key advantage of telerehabilitation is the ability to provide rehabilitation to a broad population, particularly benefiting patients in remote areas with limited access to specialized care. However, some studies report higher satisfaction and motivation in clinic-based therapy due to increased social interaction [97]. To mitigate this, approaches such as online therapist-led training [98] and multi-user interactions [99] have been proposed. Videoconferencing and family has also been found to enhance motivation and clinical outcomes by fostering accountability in patients [100].
Integrating RHHI into telerehabilitation could further enhance engagement by facilitating real-time therapist-patient physical interaction in remote settings (e.g., therapist in the clinic, patient at home). Therapists can initiate remote training sessions with a real-time assessment of the patient’s performance through remote haptic interaction paired with videoconferencing [101]. Notably, since January 2024, Medicare insurance has supported the cost of personal exoskeletons, paving the way for RHHI integration into mainstream care [102], allowing continuous rehabilitation support even after discharge.
@@ VI-A Robot Design
To date, no work has explored hand exoskeletons to establish haptic connections between individuals. Dexterous manipulation tasks, with their complexity and variability, are well-suited for RHHI, enabling users to collaboratively develop manipulation strategies. Incorporating sensory feedback, such as tactile cues about the manipulated object, could enhance interaction and promote embodiment [103, 104].
@@
RHHI for lower-limb NR should move toward multi-DoF, full-body interactions to better support complex gait and balance tasks. Recent advances have expanded RHHI to lower-limb applications, ranging from single-DoF ankle movements to more complex gait scenarios using exoskeletons or robotic manipulators. However, current implementations are limited by technological constraints. Preliminary studies on ankle tracking [29] and dyadic exoskeleton therapy [105, 106] have focused exclusively on sagittal plane actuation, neglecting essential coronal and transverse plane movements utilized in human motion. One approach to address this limitation involves end-effector robots [85] to mediate lower-limb interactions between therapists and patients. Alternatively, cable-based soft exoskeletons, [107] could facilitate dynamic interactions, though this remains largely unexplored. While both strategies improve joint alignment and synchronized movement, they may not provide sufficient power or multi-DoF assistance for more severely impaired patients. Future research should prioritize multi-axis RHHI capable of enabling comprehensive lower-limb interactions between users.
@@
RHHI must address cost and accessibility challenges without compromising therapeutic effectiveness. A major barrier to the widespread adoption of RHHI in NR is the need for multiple robotic devices, making the infrastructure prohibitively expensive for many clinics and hospitals. This financial constraint limits accessibility and practical implementation in routine therapy. To mitigate costs, Amato et al. [53] proposed IMU sensors to replace robotic devices, reducing potential costs at the expense of limiting haptic feedback. Similarly, Koh et al. [85] introduced joystick-based control, which reduces costs but can diminish the sense of embodiment. Although these approaches offer more affordable options, they come with trade-offs that could impact therapeutic effectiveness. Future research should evaluate whether these cost-saving measures adequately compensate for the loss of haptic interaction and embodiment, or if alternative approaches can be developed to balance affordability with functional benefits.
@@ VI-B Robot Control
Accurate haptic rendering through robot transparency is a fundamental requirement for effective RHHI [108]. Transparency ensures that the forces and motions experienced by the user reflect the intended interaction—minimizing the forces perceived due to the robot’s own dynamics—making it a prerequisite for delivering precise, natural, and therapeutically meaningful feedback [109, 110]. This is especially critical in NR contexts, where the fidelity of physical interaction directly affects motor learning and recovery outcomes [111, 112].
@@
Despite its importance, achieving transparency remains challenging. Most rehabilitation robots are designed for robustness and high output forces, which inherently introduces unwanted dynamics such as friction, inertia, and gravitational loads [113, 114, 115]. These effects can mask or alter the forces exchanged between human partners and robots, thereby reducing the clarity and utility of haptic cues [116, 117]. Feedforward compensation, often used to eliminate known dynamics such as gravity and friction, has been widely adopted [118, 111], but modeling inaccuracies and user variability often lead to residual errors that reduce transparency and consistency across users.
@@
RHHI effectiveness hinges not only on accurate force rendering, but also on intuitive role assignment and shared control between a user and their robot [112, 119, 111]. In this framework, the robot should function as a seamless facilitator, supporting natural, bidirectional exchange and allowing therapists to transfer their expertise directly through physical interaction. One concrete example of this is adaptive arm weight compensation [118, 120], which virtually unloads the limb to allow effortless movement while preserving haptic cues, thus enhancing both comfort and value of perceiving haptic information. Ultimately, the challenge in RHHI control lies in delivering haptic rendering that is not only technically precise, but also contextually meaningful, transforming the robot from a source of interference into a channel for physical communication [121, 122, 123].
@@
Ensuring user safety during RHHI is equally essential. While RHHI increases adaptability in therapy, it also introduces unpredictability—for example, unintentional therapist movements or patient reflexes. To mitigate these risks, Amato et al. [53] proposed filtering high-frequency components of the interaction to maintain controlled and stable coupling. Hierarchical constraint-based controllers have also been developed to prioritize safety over rendering performance when necessary [124]. Moreover, recent advances in real-time perception have enabled safe interaction with dynamic and deformable human body surfaces [125]. Intelligent systems based on machine learning and advanced collision detection can further improve safety by distinguishing intentional actions from unintended contact or perturbations [126, 127, 128].
@@ VI-C Device Networking
Ensuring patient data privacy is paramount, especially in healthcare settings where sensitive information is routinely processed. Robust encryption methods [129] are essential to protect data during transmission and prevent unauthorized access. Furthermore, strong authentication protocols must verify users’ identities, ensuring that only authorized personnel can access patient information. Compliance with regulatory frameworks [130] is also necessary to maintain high standards of data protection and confidentiality.
Unstable networks like WiFi can disrupt RHHI by causing delays, underscoring the need for reliable connectivity and advanced control. Switching to less stable networks, such as WiFi, introduces potential delays and synchronization issues, which are critical for effective bidirectional interaction in RHHI. As observed by Kim et al. [101], delays up to 500 ms were manageable for slow, controlled movements but negatively impacted fast, dynamic tasks. In fact, their communication protocol with checksum verification maintained safety for slow movements but was inadequate for fast tasks such as spasticity assessments. Similarly, Ivanova et al. [131] found that while humans can adapt to delays up to 60 ms, longer delays degrade interaction quality in the context of RHHI.
To improve performance, future research should focus on reducing latency through optimized data compression, faster communication protocols, and prioritization of critical data packets. Predictive algorithms, such as those explored by Penco et al. [132], can compensate for network delays by adjusting movement timing to maintain synchronization. Additionally, integrating local processing capabilities into RHHI devices can enable critical functions to be managed independently of network conditions, minimizing the impact of communication delays on performance.
@@ VI-D Translation of RHHI into Clinical Care
Structured and realistic training methods are essential for preparing therapists to adopt RHHI, ensuring both confidence and competence in clinical use. Studies have shown that increased training improves trust and acceptance of new rehabilitation technologies [52, 1, 133]. This is particularly important in RHHI, where therapists are not only facilitators but also end-users alongside patients. For this purpose, therapist-standardized procedures should be implemented. For instance, in Kucuktabak et al. [134], therapists participated in training sessions where they were haptically connected to healthy team members simulating asymmetric gait before working with patients. An alternative solution is to artificially induce impairments in healthy participants to simulate neuromotor deficits. Manczurowsky et al. [135] introduced ”dysfunctional electrical stimulation” to involuntarily activate the hamstrings of healthy participants during walking, mimicking neuromotor impairments. This method could improve therapist training by exposing them to realistic patient challenges in a controlled environment [136]. Hasson et al. [137, 138] developed a system where participants modified a virtual stroke survivor’s gait through robotic manipulanda. Simulated training methods provide therapists with unlimited exposure to patient movement dynamics, reducing risks associated with real patient scenarios. Additionally, impairment simulations using visual noise or mechanical constraints [139, 140] can further enhance training realism.
@@
RHHI should support a broad and functional range of exercises through standardized yet customizable task libraries to meet diverse NR needs. A key advantage of RHHI over conventional robotic rehabilitation is its ability to support a diverse range of exercises. Upper-limb RHHI studies primarily focus on stereotyped reaching or tracking tasks [25], which lack functional relevance for daily activities. Future development should incorporate more naturalistic tasks, such as grasping and object manipulation, to improve real-world applicability [1]. For lower-limb rehabilitation, RHHI applications range from single-joint tracking tasks targeting isolated motor control [29], to multi-joint functional exercises, such as treadmill walking, sit-to-stand transitions, and obstacle negotiation [53]. A standardized exercise library should be developed for RHHI platforms while allowing therapists to customize activities for individual patient needs. This balance between standardization and adaptability will ensure that RHHI systems effectively address diverse rehabilitation goals.
Box 2: Learning by Demonstration and RHHI What is it? Learning by Demonstration (LbD) models therapist strategies from dyadic interactions, enabling adaptive controllers that improve robotic assistance and patient outcomes [141]. Why LbD and RHHI? (1) Captures therapist-patient interaction forces during physical guidance; (2) Enhances therapist trust and promotes patient-specific adaptation; (3) Enables robots to learn assistance strategies without needing hard-coded rules and improve over time through ongoing feedback and demonstration. Recent examples: Upper-limb exoskeleton - Luciani et al. [142, 55] introduced an LbD framework that captures therapist interaction forces while guiding a patient’s movement using an upper-limb exoskeleton. Lower-limb exoskeleton - Sankar et al. [143] demonstrated preliminary results in this direction, using an LbD model trained with data from a therapist-patient dyad connected to lower-limb exoskeletons. What is next? Incorporating additional sensor data (e.g., heart rate, skin conductance) could further refine control policies. These adaptive systems could improve over time through ongoing therapist feedback and demonstrations. Takeaway: Applying LbD to RHHI can offer insights into therapist-patient interactions, enabling the design of models that learn personalized assistive or resistive strategies to the patient’s needs.
@@ VII Open research directions
In this section, we outline key open research questions in the field of RHHI. We believe that addressing our raised questions will help identify the impact of this technology on NR, the factors that contribute to its effectiveness, and maximize its potential for clinical implementation.
@@ VII-A What are the foundational mechanisms behind Motor Learning in RHHI in NR?
Numerous studies on RHHI in healthy individuals have explored gains in individual performance (i.e., motor learning) following haptic training with a partner [144]. These studies often use trajectory tracking tasks, sometimes incorporating perturbations (e.g., force fields, visuomotor rotations) to examine changes in task performance. Some findings suggest that haptic training with a partner accelerates motor learning compared to training alone [23, 145] or compared to training with a direct connection to the target trajectory [46]. While gains have been attributed to mutual information sharing through the haptic connection [73, 24], these learning benefits are not consistently observed across all studies.
Research has found no significant differences in individual learning outcomes between haptically connected and unconnected training in certain contexts, particularly during multi-joint reaching tasks [146] and lower-limb ankle tracking tasks [147]. These findings suggest that, in such cases, improvements may result more from the mechanics of interaction—such as error averaging across users—than from mutual adaptation of movement strategies, which is more commonly observed in upper-limb studies [29]. These heterogeneous results highlight that the learning effects of RHHI are likely influenced by multiple factors, including robotic system characteristics (e.g., quality of haptic rendering), task-specific constraints (e.g., DoFs, workspace size, and difficulty), and the limbs involved (e.g. upper or lower ones).
@@
Given these differences, the mechanisms of motor learning in RHHI for NR are likely to be more complex than those observed in controlled laboratory settings with healthy individuals [73, 24]. Unlike experimental studies that isolate haptic input, clinical applications will involve the combined influence of multisensory feedback and cognitive engagement. While prior work provides a useful foundation, further research is needed to characterize how learning unfolds when patients and therapists interact through RHHI with full awareness of the physical and social dynamics involved.

### v2

@@ V-F Remote Therapy
Telerehabilitation (rehabilitation delivered remotely) has proven effective in improving motor function, cognition, balance, speech, and communication skills [nikolaev2022recent]. Studies have shown that telerehabilitation can yield outcomes comparable, or even superior, to conventional rehabilitation [pitliya2025telerehabilitation]. A key advantage of telerehabilitation is the ability to provide rehabilitation to a broad population, particularly benefiting patients in remote areas with limited access to specialized care. However, some studies report higher satisfaction and motivation in clinic-based therapy due to increased social interaction [cramer2019efficacy]. To mitigate this, approaches such as online therapist-led training [smith2020combined] and multi-user interactions [thielbar2020home] have been proposed. Videoconferencing and family has also been found to enhance motivation and clinical outcomes by fostering accountability in patients [chen2020qualitative]. RHHI can further support rehabilitation when physical proximity is limited, such as for patients requiring social distancing or individuals with Autism. In these cases, robots can act as mediators that facilitate engagement and communication between the patient and therapist, promoting social interaction, emotional expression, and even motor training [santos2021design, fachantidis2020tauhe, jouaiti2019robot].
RHHI can overcome telerehabilitation physical distance constraints [nuara2022telerehabilitation], enhancing engagement by facilitating real-time therapist-patient physical interaction in remote settings (e.g., therapist in the clinic, patient at home). Therapists can initiate remote training sessions with a real-time assessment of the patient’s performance through remote haptic interaction paired with videoconferencing [kim2020remote]. Alternatively, patients could connect to each other or to remote family members over the Internet and perform competitive, collaborative, or cooperative exercises together while tele/videoconferencing with each other, similarly to how modern videogames offer online play [thielbar2020home]. Notably, since January 2024, Medicare insurance has supported the cost of personal exoskeletons, paving the way for RHHI integration into mainstream care [ExoMedicaid], allowing continuous rehabilitation support even after discharge.
@@ VI-A Robot Design
To date, no work has explored hand exoskeletons to establish haptic connections between individuals. Dexterous manipulation tasks, with their complexity and variability, are well-suited for RHHI, enabling users to collaboratively develop manipulation strategies. Incorporating sensory feedback, such as haptic cues about the manipulated object, could enhance interaction and promote embodiment [pinardi2023impact, noccaro4884432vibrotactile].
@@
RHHI for lower-limb NR should move toward multi-DoF, full-body interactions to better support complex gait and balance tasks. Recent advances have expanded RHHI to lower-limb applications, ranging from single-DoF ankle movements to more complex gait scenarios using exoskeletons or robotic manipulators. However, current implementations are limited by technological constraints. Preliminary studies on ankle tracking [short2023haptic] and dyadic exoskeleton therapy [vianello2023exoskeleton, kuccuktabak2023virtual] have focused exclusively on sagittal plane actuation, neglecting essential coronal and transverse plane movements utilized in human motion. One approach to address this limitation involves end-effector robots [koh2021exploiting] to mediate lower-limb interactions between therapists and patients. Alternatively, cable-based soft exoskeletons, [basla2024enhancing] could facilitate dynamic interactions, though this remains largely unexplored. While both strategies improve joint alignment and synchronized movement, they may not provide sufficient power or multi-DoF assistance for more severely impaired patients. Future research should prioritize multi-axis RHHI capable of enabling comprehensive lower-limb interactions between users.
@@
RHHI must address cost and accessibility challenges without compromising therapeutic effectiveness. A major barrier to the widespread adoption of RHHI in NR is the need for multiple robotic devices, making the infrastructure prohibitively expensive for many clinics and hospitals. This financial constraint limits accessibility and practical implementation in routine therapy. To mitigate costs, Amato et al. [amato2024unidirectional] proposed IMU sensors to replace robotic devices, reducing potential costs at the expense of limiting haptic feedback. Similarly, Koh et al. [koh2021exploiting] introduced joystick-based control, which reduces costs but can diminish the sense of embodiment. Although these approaches offer more affordable options, they come with trade-offs that could impact therapeutic effectiveness. Future research should evaluate whether these cost-saving measures adequately compensate for the loss of haptic interaction and embodiment, or if alternative approaches can be developed to balance affordability with functional benefits. At the same time, design under co-creation schemes involving clinicians and engineers ensures patient engagement and adherence to therapies [dobe2023co].
Box 2: Learning by Demonstration and RHHI What is it? Learning by Demonstration (LbD) models therapist strategies from dyadic interactions, enabling adaptive controllers that improve robotic assistance and patient outcomes [mulian2024mimicking]. Why LbD and RHHI? (1) Captures therapist-patient interaction forces during physical guidance; (2) Enhances therapist trust and promotes patient-specific adaptation; (3) Enables robots to learn assistance strategies without needing hard-coded rules and improve over time through ongoing feedback and demonstration. Recent examples: (1) Upper-limb exoskeleton - Luciani et al. [luciani2024imitation, luciani2024therapists] introduced an LbD framework that captures therapist interaction forces while guiding a patient’s movement using an upper-limb exoskeleton. (2) Lower-limb exoskeleton - Sankar et al. [sankar2024action] demonstrated preliminary results in this direction, using an LbD model trained with data from a therapist-patient dyad connected to lower-limb exoskeletons. What is next? Incorporating additional sensor data (e.g., heart rate, skin conductance) could further refine control policies. These adaptive systems could improve over time through ongoing therapist feedback and demonstrations. Takeaway: Applying LbD to RHHI can offer insights into therapist-patient interactions, enabling the design of models that learn personalized assistive or resistive strategies to the patient’s needs.
@@ VI-B Robot Control
Accurate haptic rendering through robot transparency is a fundamental requirement for effective RHHI [kuccuktabak2023haptic]. Transparency ensures that the forces and motions experienced by the user reflect the intended interaction—minimizing the forces perceived due to the robot’s own dynamics—making it a prerequisite for delivering precise, natural, and therapeutically meaningful feedback [lawrence1993stability, colgate1997passivity]. This is especially critical in NR contexts, where the fidelity of physical interaction directly affects motor learning and recovery outcomes [proietti2016upper, jarrasse2014perspective].
@@
Despite its importance, achieving transparency remains challenging. Most rehabilitation robots are designed for robustness and high output forces, which inherently introduces unwanted dynamics such as friction, inertia, and gravitational loads [hogan1985impedance, just2016motor, just2017feedforward]. These effects can mask or alter the forces exchanged between human partners and robots, thereby reducing the clarity and utility of haptic cues [noccaro2023evaluation, just2018exoskeleton]. Feedforward compensation, often used to eliminate known dynamics such as gravity and friction, has been widely adopted [jarrasse2011connecting, proietti2016upper], but modeling inaccuracies and user variability often lead to residual errors that reduce transparency and consistency across users.
@@
RHHI effectiveness hinges not only on accurate force rendering, but also on intuitive role assignment and shared control between a user and their robot [jarrasse2014perspective, jarrasse2014slaves, proietti2016upper]. In this framework, the robot should function as a seamless facilitator, supporting natural, bidirectional exchange and allowing therapists to transfer their expertise directly through physical interaction. One concrete example of this is adaptive arm weight compensation [jarrasse2011connecting, just2020human], which virtually unloads the limb to allow effortless movement while preserving haptic cues, thus enhancing both comfort and value of perceiving haptic information. Ultimately, the challenge in RHHI control lies in delivering haptic rendering that is not only technically precise, but also contextually meaningful, transforming the robot from a source of interference into a channel for physical communication [reinkensmeyer2004robotics, krebs1998robot, ajoudani2018progress].
@@
Ensuring user safety during RHHI is equally essential. While RHHI increases adaptability in therapy, it also introduces unpredictability—for example, unintentional therapist movements or patient reflexes. To mitigate these risks, Amato et al. [amato2024unidirectional] proposed filtering high-frequency components of the interaction to maintain controlled and stable coupling. Hierarchical constraint-based controllers have also been developed to prioritize safety over rendering performance when necessary [kuccuktabak2024physical]. Moreover, recent advances in real-time perception have enabled safe interaction with dynamic and deformable human body surfaces [Sommerhalder]. Intelligent systems based on machine learning and advanced collision detection can further improve safety by distinguishing intentional actions from unintended contact or perturbations [haddadin2017robot, deluca2006collision, haddadin2017robotcollisions].
@@ VI-C Device Networking
Ensuring patient data privacy is paramount, especially in healthcare settings where sensitive information is routinely processed. Robust encryption methods [deng2022security] are essential to protect data during transmission and prevent unauthorized access. Furthermore, strong authentication protocols must verify users’ identities, ensuring that only authorized personnel can access patient information. Compliance with regulatory frameworks [maddali2024ai] is also necessary to maintain high standards of data protection and confidentiality.
Unstable networks like WiFi can disrupt RHHI by causing delays, underscoring the need for reliable connectivity and advanced control. Switching to less stable networks introduces potential delays and synchronization issues, which are critical for effective bidirectional interaction in RHHI. As observed by Kim et al. [kim2020remote], delays up to 500 ms were manageable for slow, controlled movements but negatively impacted fast, dynamic tasks. In fact, their communication protocol with checksum verification maintained safety for slow movements but was inadequate for fast tasks such as spasticity assessments. Similarly, Ivanova et al. [ivanova2021short] found that while humans can adapt to delays up to 60 ms, longer delays degrade interaction quality in the context of RHHI.
To improve performance, future research should focus on reducing latency through optimized data compression, faster communication protocols, and prioritization of critical data packets. Predictive algorithms, such as those explored by Penco et al. [penco2023prescient], can compensate for network delays by adjusting movement timing to maintain synchronization. Additionally, integrating local processing capabilities into RHHI devices can enable critical functions to be managed independently of network conditions, minimizing the impact of communication delays on performance.
@@ VI-D Translation of RHHI into Clinical Care
Structured and realistic training methods are essential for preparing therapists to adopt RHHI, ensuring both confidence and competence in clinical use. Studies have shown that increased training improves trust and acceptance of new rehabilitation technologies [hasson2023neurorehabilitation, celian2021day, mortenson2022therapists]. This is particularly important in RHHI, where therapists are not only facilitators but also end-users alongside patients. For this purpose, therapist-standardized procedures should be implemented. For instance, in Kucuktabak et al. [kucuktabak2025therapist], therapists participated in training sessions where they were haptically connected to healthy team members simulating asymmetric gait before working with patients. An alternative solution is to artificially induce impairments in healthy participants to simulate neuromotor deficits. Manczurowsky et al. [manczurowsky2024induced] introduced ”dysfunctional electrical stimulation” to involuntarily activate the hamstrings of healthy participants during walking, mimicking neuromotor impairments. This method could improve therapist training by exposing them to realistic patient challenges in a controlled environment [buddeoutside]. Hasson et al. [hasson2019learning, hasson2019visual] developed a system where participants modified a virtual stroke survivor’s gait through robotic manipulanda. Simulated training methods provide therapists with unlimited exposure to patient movement dynamics, reducing risks associated with real patient scenarios. Additionally, impairment simulations using visual noise or mechanical constraints [noccaro2022visual, buscaglione2023assessment] can further enhance training realism.
@@
RHHI should support a broad and functional range of exercises through standardized yet customizable task libraries to meet diverse NR needs. A key advantage of RHHI over conventional robotic rehabilitation lies in its ability to support a broader range of exercises. However, most existing RHHI studies for both upper and lower limbs focus on stereotyped reaching or tracking tasks [kuccuktabak2021human], which have limited functional relevance. Future developments should incorporate more naturalistic and multi-joint functional tasks, e.g., object manipulation, and obstacle negotiation [celian2021day, amato2024unidirectional]. Establishing a standardized yet customizable exercise library, developed in collaboration with therapists, would balance consistency and adaptability to effectively address diverse rehabilitation goals. In Box. VI-D, a Road Map to Clinical Adoption is summarized.
Box 3: A Road Map to Clinical Adoption Co-Creation and Stakeholder Involvement - The success of robot-mediated Human–Human Interaction (RHHI) technologies relies on co-creation among patients, clinicians, engineers, and caregivers to ensure clinical relevance, usability, and robustness. Therapists should be engaged early through workshops and iterative design feedback to align features with therapeutic needs [nicora2025healthcare, aprile2025rehabilitation]. Group discussions with therapists are valuable for identifying barriers and expectations, as shown in prior Technology Acceptance Model studies [luciani2023technology]. Effective clinical translation of RHHI requires coordinated efforts across engineering, clinical, regulatory, and institutional domains—addressing training, certification, workflow alignment, and safety—while relying on therapists as key mediators who integrate these technologies into practice and bridge engineering design, patient needs, and healthcare delivery [thawisuk2025perspectives, huang2018translation]. Adaptation and Clinical Integration - Effective clinical adoption of RHHI requires structured training and defined competency pathways, including didactic instruction, supervised practice, standardized evaluations, and periodic re-certification. Integration into existing workflows should be supported by pilot studies assessing feasibility, therapist acceptance, and patient safety, with clearly delineated staff roles and safety protocols. Embedding RHHI within established frameworks such as the International Classification of Functioning, Disability and Health, and aligning it with clinical assessment scales will facilitate implementation. Practical considerations are critical: therapists value minimal setup time, reliable operation, and compatibility with typical 45–60-minute sessions. Legal and Regulatory Considerations - Compliance with regulatory and ethical standards (HIPAA, GDPR, IEC/ISO, FDA/MDR) is essential, alongside clear attention to liability, cost, and maintenance. Embedding safety, usability, and risk management from the design stage—supported by post-market monitoring—ensures RHHI systems are reliable, accepted, and ready for routine clinical use[ali2025ethical].
@@ VII Open research directions
In this section, we outline key open research questions in the field of RHHI that will help identify the impact of this technology on NR, the factors that contribute to its effectiveness, and maximize its potential for clinical implementation.
@@ VII-A What are the foundational mechanisms behind motor learning in RHHI in NR?
Numerous studies on RHHI in healthy individuals have explored gains in individual performance (i.e., motor learning) following haptic training with a partner [waters2024motor]. These studies often use trajectory tracking tasks, sometimes incorporating perturbations (e.g., force fields, visuomotor rotations) to examine changes in task performance. Some findings suggest that haptic training with a partner accelerates motor learning compared to training alone [Ganesh2014, Batson2020] or compared to training with a direct connection to the target trajectory [ivanova2022interaction]. While gains have been attributed to mutual information sharing through the haptic connection [takagi2017physically, Takagi2018HapticInteraction], these learning benefits are not consistently observed across all studies.
Research has found no significant differences in individual learning outcomes between haptically connected and unconnected training in certain contexts, particularly during multi-joint reaching tasks [Beckers2020] and lower-limb ankle tracking tasks [kim2022effect]. These findings suggest that, in such cases, improvements may result more from the mechanics of interaction—such as error averaging across users—than from mutual adaptation of movement strategies, which is more commonly observed in upper-limb studies [short2023haptic]. These heterogeneous results highlight that the learning effects of RHHI are likely influenced by multiple factors, including robotic system characteristics (e.g., quality of haptic rendering), task-specific constraints (e.g., DoFs, workspace size, and difficulty), and the limbs involved (e.g. upper or lower ones).
@@
Given these differences, the mechanisms of motor learning in RHHI for NR are likely to be more complex than those observed in controlled laboratory settings with healthy individuals [takagi2017physically, Takagi2018HapticInteraction]. Unlike experimental studies that isolate haptic input, clinical applications will involve the combined influence of multisensory feedback and cognitive engagement. For example, a recent study that tried to combine haptic interaction with engaging visual feedback failed to replicate previously observed motor learning benefits as a result of haptic interaction between partners [hossain2025combining]. While prior work provides a useful foundation, further research is needed to characterize how learning unfolds when patients and therapists interact through RHHI with full awareness of the physical and social dynamics involved.